# DQ-LoRe: Dual Queries with Low Rank Approximation Re-ranking for In-Context Learning

### v5

@@ Appendix G Case Study on SVAMP
Question: Mary is baking a cake. The recipe calls for 14 cups of flour and 6 cups of sugar. She already put in 5 cups of flour.How many more cups of flour than cups of sugar does she need to add now?
A: Let’s think step by step Mary needs to add 14 - 5 = 9 more cups of flour. She needs to add 9 - 6 = 3 more cups of flour than cups of sugar. Answer: 333. The answer is 3.
Question: Mary is baking a cake. The recipe calls for 11 cups of flour and 7 cups of sugar. She already put in some cups of flour. If she still needs 2 more cups of flour than sugarHow many cups of flour did she put in?
A: Let’s think step by step. Let’s assume Mary already put x cups of flour. According to the information given, she still needs 2 more cups of flour than sugar, so she needs (7 + 2) = 9 cups of sugar. Since the recipe calls for 11 cups of flour, the total amount of flour needed is 11 cups. Therefore, the equation is x + 9 = 11. Subtracting 9 from both sides of the equation, we get x = 11 - 9 = 2. Mary has already put in 2 cups of flour. The answer is 2.
Question: Mary is baking a cake. The recipe calls for 11 cups of sugar and some more cups of flour. She already put in 3 cups of flour. If she still needs to add 6 more cups of flourHow many cups of flour did the recipe require?
A: Let’s think step by step: Mary already put in 3 cups of flour, and she still needs to add 6 more cups of flour. So, the total amount of flour she needs to add is 3 + 6 = 9 cups. Therefore, the recipe requires 9 cups of flour. The answer is 9.
Question: Mary is baking a cake. The recipe calls for 3 cups of sugar 10 cups of flour and 15 cups of salt. She already put in 6 cups of flour.How many more cups of flour does she need to add?
A: Let’s think step by step To find out how many more cups of flour Mary needs to add, we need to subtract the amount of flour she already put in from the total amount of flour required. The total amount of flour required is 10 cups, and Mary already put in 6 cups. Therefore, she needs to add 10 - 6 = 4 more cups of flour. Mary needs to add 4 cups of flour. The answer is 4.
Table 13: EPR Case Study on SVAMP: Part 2
  Problem Question: Mary is baking a cake. The recipe calls for 9 cups of flour and 11 cups of sugar. She already put in 4 cups of flour.How many more cups of sugar than cups of flour does she need to add now?
A: Let’s think step by step Mary needs to add 9 - 4 = 5 more cups of flour. Mary needs to add 11 - 5 = 6 more cups of sugar than cups of flour. Answer: 666. The answer is 6.
Question: Mary is baking a cake. The recipe calls for 12 cups of sugar and 14 cups of flour. She already put in 10 cups of sugar.How many more cups of flour than cups of sugar does she need to add now?
A: Let’s think step by step The recipe calls for 12 cups of sugar and she already put in 10 cups, so she needs to add 12 - 10 = 2 more cups of sugar. The recipe calls for 14 cups of flour and she hasn’t added any yet, so she needs to add 14 cups of flour. To find the difference between the cups of flour and cups of sugar, we subtract the cups of sugar from the cups of flour: 14 - 2 = 12. She needs to add 12 more cups of flour than cups of sugar. The answer is 12.
Table 14: CEIL Case Study on SVAMP
  Problem
Question: Mary is baking a cake. The recipe calls for 6 cups of flour 8 cups of sugar and 7 cups of salt. She already put in 5 cups of flour.How many more cups of sugar than cups of salt does she need to add now?
CEIL Prompt
Question: A mailman has to give 4 pieces of junk mail to each house in each of the 81 blocks. If there are 12 houses in each blockHow many pieces of junk mail should he give in each block?
A: Let’s think step by step There are 12 houses in each block, and the mailman has to give 4 pieces of junk mail to each house. So, in each block, the mailman should give 12 houses x 4 pieces of junk mail = 48 pieces of junk mail. The answer is 48.
Question: After eating a hearty meal they went to see the Buckingham palace. There, Rachel learned that 132 visitors came to the Buckingham palace that day. If 406 people visited the Buckingham palace within the past 327 daysHow many visitors visited the Buckingham palace on the previous day?
A: Let’s think step by step: To find out how many visitors visited the Buckingham palace on the previous day, we need to subtract the number of visitors on the current day from the total number of visitors in the past 327 days. So, the number of visitors on the previous day is 406 - 132 = 274. Therefore, 274 visitors visited the Buckingham palace on the previous day. The answer is 274.
Question: There are a total of 16 peaches in a basket. If there are 13 red peaches and some more green peaches in the basket.How many green peaches are in the basket?
A: Let’s think step by step If there are a total of 16 peaches in the basket and 13 of them are red, then the remaining peaches must be green. So, the number of green peaches in the basket is 16 - 13 = 3. The answer is 3.
Question: If each bag has 41 cookies and you had 53 bags of cookiesHow many cookies would you have?
A: Let’s think step by step If each bag has 41 cookies, and you have 53 bags, then the total number of cookies would be 41 x 53 = 2173 cookies. The answer is 2173.
Question: The Ferris wheel in paradise park has 2 small seats and 23 large seats. Each small seat can hold 14 people and large seat can hold 54 people.How many people can ride the Ferris wheel on small seats?
A: Let’s think step by step The number of people that can ride on the small seats is equal to the number of small seats multiplied by the capacity of each small seat. There are 2 small seats and each small seat can hold 14 people, so the total number of people that can ride on the small seats is 2 x 14 = 28. The answer is 28.
Question: Frank was reading through his favorite book. The book had 41 chapters, each with the same number of pages. It has a total of 450 pages. It took Frank 30 days to finish the book.How many pages did he read per day?
A: Let’s think step by step. If the book has a total of 450 pages and 41 chapters, then each chapter has 450/41 = 10.975609756097561 pages (rounded to the nearest whole number). Since each chapter has the same number of pages, Frank read approximately 11 pages per chapter. If Frank finished the book in 30 days, then he read 450/30 = 15 pages per day. Therefore, Frank read approximately 15 pages per day. The answer is 15.
Question: Julia played tag with 16 kids on monday. If she played tag with 12 more kids on monday than on tuesdayHow many kids did she play with on tuesday?
A: Let’s think step by step: If Julia played with 16 kids on Monday and 12 more kids than on Tuesday, then the number of kids she played with on Tuesday is 16 - 12 = 4. Therefore, Julia played with 4 kids on Tuesday. The answer is 4.
Question: Lewis earns \$2 every week during the harvest. If he earns a total of \$ 178How many weeks did the harvest last?
A: Let’s think step by step: If Lewis earns \$2 every week during the harvest, and he earns a total of \$178, we can divide the total earnings by the amount he earns per week to find the number of weeks. \$178 / \$2 = 89 weeks Therefore, the harvest lasted for 89 weeks. The answer is 89.
Table 15: DQ-LoRe Case Study on SVAMP: Part 1
  Problem
Question: Mary is baking a cake. The recipe calls for 6 cups of flour 8 cups of sugar and 7 cups of salt. She already put in 5 cups of flour.How many more cups of sugar than cups of salt does she need to add now?
DQ-LoRe (ours)
Question: Mary is baking a cake. The recipe calls for 5 cups of sugar and 14 cups of flour. She already put in 11 cups of flour.How many more cups of sugar than cups of flour does she need to add now?
A: Let’s think step by step. The recipe requires 5 cups of sugar and 14 cups of flour.She already added 11 cups of flour, so she still needs to add 14 - 11 = 3 cups of flour.For the sugar, she needs 5 cups and she hasn’t added any yet, so she needs to add all 5 cups.Therefore, she needs to add 5 - 3 = 2 more cups of sugar than cups of flour. The answer is 2.
Question: Mary is baking a cake. The recipe calls for 9 cups of sugar 7 cups of flour and 4 cups of salt. She already put in 2 cups of flour.How many more cups of flour than cups of salt does she need to add now?
A: Let’s think step by step. Mary needs to add 7 - 2 = 5 more cups of flour. She needs to add 5 - 4 = 1 more cup of flour than cups of salt. Answer: 111. The answer is 1.
Question: Mary is baking a cake. The recipe calls for 11 cups of sugar and some more cups of flour. She already put in 3 cups of flour. If she still needs to add 6 more cups of flour.How many cups of flour did the recipe require?
A: Let’s think step by step. Mary already put in 3 cups of flour, and she still needs to add 6 more cups of flour. So, the total amount of flour she needs to add is 3 + 6 = <<3+6=9>>9 cups. Therefore, the recipe requires 9 cups of flour. The answer is 9.
Question: Mary is baking a cake. The recipe calls for 6 cups of sugar and 14 cups of flour. She already put in 7 cups of flour and 60 cups of sugar.How many more cups of flour does she need to add?
A: Let’s think step by step. To find out how many more cups of flour Mary needs to add, we need to subtract the amount of flour she already put in from the total amount required. Total cups of flour required = 14 cups Flour already put in = 7 cups To find the remaining amount of flour needed, we do: Remaining flour needed = Total flour required - Flour already put in Remaining flour needed = 14 cups - 7 cups Remaining flour needed = 7 cups Mary needs to add 7 more cups of flour. The answer is 7.
Question: Helen the hippo and her friends are preparing for thanksgiving at Helen’s house. Helen baked 519 chocolate chip cookies and 300 raisin cookies yesterday. And she baked 280 raisin cookies and 359 chocolate chip cookies this morning.How many more raisin cookies did Helen bake yesterday compared to today?
A: Let’s think step by step. To find out how many more raisin cookies Helen baked yesterday compared to today, we need to subtract the number of raisin cookies she baked today from the number she baked yesterday. Yesterday, Helen baked 519 chocolate chip cookies and 300 raisin cookies, for a total of 519 + 300 = 819 cookies. Today, Helen baked 280 raisin cookies and 359 chocolate chip cookies, for a total of 280 + 359 = 639 cookies. To find the difference in the number of raisin cookies, we subtract the number of raisin cookies baked today from the number baked yesterday: 300 - 280 = 20. Therefore, Helen baked 20 more raisin cookies yesterday compared to today. The answer is 20.
Question: Mary is baking a cake. The recipe calls for 5 cups of flour and 14 cups of sugar. She already put in 10 cups of flour and 2 cups of sugar.How many more cups of sugar does she need to add?
A: Let’s think step by step. To find out how many more cups of sugar Mary needs to add, we need to subtract the amount of sugar she already put in from the total amount required. The recipe calls for 14 cups of sugar and Mary already put in 2 cups, so she still needs to add 14 - 2 = 12 more cups of sugar. Therefore, Mary needs to add 12 more cups of sugar. The answer is 12.
The answer is 2.
Table 16: DQ-LoRe Case Study on SVAMP: Part 2
  Problem Question: Matthew gave equal numbers of crackers and cakes to his 4 friends. If he had 32 crackers and 98 cakes initially. How many crackers did each person eat?
A: Let’s think step by step. If Matthew gave equal numbers of crackers and cakes to his 4 friends, then each friend received the same number of crackers and cakes. Let’s assume that each friend received x number of crackers and x number of cakes. Since Matthew initially had 32 crackers and he gave an equal number to each friend, the total number of crackers given to his friends is 4x. Similarly, since Matthew initially had 98 cakes and he gave an equal number to each friend, the total number of cakes given to his friends is also 4x. We can set up the equation: 4x = 32 Solving for x, we divide both sides of the equation by 4: x = 32/4 = 8 Therefore, each person ate 8 crackers. The answer is 8.
Question: Rachel had to complete 5 pages of math homework and yet more pages of reading homework. If she had to complete a total of 7 pages of homework. How many pages of reading homework did she have to complete?
A: Let’s think step by step. Rachel had to complete 5 pages of math homework and yet more pages of reading homework. If she had to complete a total of 7 pages of homework, then the number of reading homework pages can be found by subtracting the number of math homework pages from the total number of homework pages. So, Rachel had to complete 7 - 5 = 2 pages of reading homework. Therefore, Rachel had to complete 2 pages of reading homework.

### v6

@@ Appendix F Case Study on SVAMP
EPR Retrieved Exemplars (Continued) More EPR Retrieved Exemplars: Exemplar 6: Mary is baking a cake. The recipe calls for 9 cups of flour and 11 cups of sugar. She already put in 4 cups of flour. How many more cups of sugar than cups of flour does she need to add now? A: Mary needs to add 9 - 4 = 5 more cups of flour. She needs to add 11 - 5 = 6 more cups of sugar than cups of flour. Exemplar 7: Mary is baking a cake. The recipe calls for 12 cups of sugar and 14 cups of flour. She already put in 10 cups of sugar. How many more cups of flour than cups of sugar does she need to add now? A: She needs to add 12 - 10 = 2 more cups of sugar, and 14 cups of flour. Difference: 14 - 2 = 12 more cups of flour than cups of sugar. Analysis: All EPR exemplars share the same “Mary baking a cake” context, demonstrating EPR’s reliance on surface-level word co-occurrence rather than reasoning patterns.
CEIL Retrieved Exemplars Observation: CEIL retrieves diverse but often irrelevant exemplars that lack connection to the test problem’s reasoning structure. Exemplar 1: A mailman has to give 4 pieces of junk mail to each house in each of the 81 blocks. If there are 12 houses in each block, how many pieces of junk mail should he give in each block? A: 12 houses × 4 pieces = 48 pieces of junk mail per block. Exemplar 2: Rachel learned that 132 visitors came to the Buckingham palace that day. If 406 people visited within the past 327 days, how many visitors visited on the previous day? A: 406 - 132 = 274 visitors on the previous day. Exemplar 3: There are a total of 16 peaches in a basket. If there are 13 red peaches, how many green peaches are in the basket? A: 16 - 13 = 3 green peaches. Exemplar 4: If each bag has 41 cookies and you had 53 bags of cookies, how many cookies would you have? A: 41 × 53 = 2173 cookies. Exemplar 5: The Ferris wheel has 2 small seats (14 people each). How many people can ride on small seats? A: 2 × 14 = 28 people. Exemplar 6: Frank’s book had 450 pages. It took 30 days to finish. How many pages did he read per day? A: 450 / 30 = 15 pages per day. Analysis: CEIL retrieves exemplars with diverse contexts (mailman, palace, peaches, cookies, Ferris wheel, book) that are semantically unrelated to the test problem about baking. This demonstrates CEIL’s focus on diversity over relevance.
DQ-LoRe Retrieved Exemplars (Ours) Observation: DQ-LoRe retrieves exemplars that share reasoning patterns with the test problem, not just surface-level word co-occurrence. Exemplar 1: Mary is baking a cake. The recipe calls for 5 cups of sugar and 14 cups of flour. She already put in 11 cups of flour. How many more cups of sugar than cups of flour does she need to add now? A: She needs 14 - 11 = 3 more cups of flour, and 5 cups of sugar. Difference: 5 - 3 = 2 more cups of sugar than flour. Exemplar 2: Mary is baking a cake. The recipe calls for 9 cups of sugar, 7 cups of flour and 4 cups of salt. She already put in 2 cups of flour. How many more cups of flour than cups of salt does she need to add now? A: She needs 7 - 2 = 5 more cups of flour. Difference: 5 - 4 = 1 more cup of flour than salt. Exemplar 3: Mary is baking a cake. The recipe calls for 6 cups of sugar and 14 cups of flour. She already put in 7 cups of flour. How many more cups of flour does she need? A: 14 - 7 = 7 more cups of flour. Exemplar 4: Helen baked 300 raisin cookies yesterday. She baked 280 raisin cookies today. How many more raisin cookies did she bake yesterday compared to today? A: 300 - 280 = 20 more raisin cookies yesterday. (Different context but same “comparison” reasoning pattern) Analysis: DQ-LoRe retrieves exemplars that share the same reasoning pattern (subtraction followed by comparison) with the test problem. Notably, Exemplar 4 has a completely different context (cookies vs. baking ingredients) but shares the same mathematical reasoning structure.
DQ-LoRe Retrieved Exemplars (Ours) (Continued) Additional DQ-LoRe Examples: Exemplar 5: Matthew gave equal numbers of crackers to his 4 friends. If he had 32 crackers initially, how many crackers did each person eat? A: 32 / 4 = 8 crackers per person. Exemplar 6: Rachel had to complete 5 pages of math homework and more pages of reading homework. If total homework was 7 pages, how many pages of reading homework? A: 7 - 5 = 2 pages of reading homework. Key Insight: DQ-LoRe’s dual query mechanism (question + CoT) enables it to find exemplars with similar reasoning structures, even when the surface-level context differs significantly. This leads to more robust in-context learning compared to methods that rely on word co-occurrence (EPR) or diversity alone (CEIL).